# {cutout}00.3cm22cm1 empty    PointMamba: A Simple State Space Model for Point Cloud Analysis

### v4

@@ Abstract
Transformers have become one of the foundational architectures in point cloud analysis tasks due to their excellent global modeling ability. However, the attention mechanism has quadratic complexity, making the design of a linear complexity method with global modeling appealing. In this paper, we propose PointMamba [Uncaptioned image], transferring the success of Mamba, a recent representative state space model (SSM), from NLP to point cloud analysis tasks. Unlike traditional Transformers, PointMamba employs a linear complexity algorithm, presenting global modeling capacity while significantly reducing computational costs. Specifically, our method leverages space-filling curves for effective point tokenization and adopts an extremely simple, non-hierarchical Mamba encoder as the backbone. Comprehensive evaluations demonstrate that PointMamba achieves superior performance across multiple datasets while significantly reducing GPU memory usage and FLOPs. This work underscores the potential of SSMs in 3D vision-related tasks and presents a simple yet effective Mamba-based baseline for future research. The code is available at https://github.com/LMD0311/PointMamba.
@@ 1 Introduction
Point cloud analysis is one of the fundamental tasks in computer vision and has a wide range of real-world applications shi2020points ; wu2023point ; chen2023voxelnext , including robotics, autonomous driving, and augmented reality. It is a challenging task due to the intrinsic irregularity and sparsity of point clouds. To address the issues, there has been rapid progress in deep learning-based methods pang2022masked ; liu2022masked ; qian2022pointnext ; zhou2024dynamic , consistently pushing the performance to the new record.
@@
Inspired by this, this paper aims to unlock the potential of SSM in point cloud analysis tasks, discussing whether it can be a viable alternative to Transformers in this domain. Through a series of pilot experiments, we find that directly using the pioneering SSM, Mamba gu2023mamba , can not achieve ideal performance. We argue that the main inherent limitation comes from the unidirectional modeling employed by the default Mamba, as the context is obtained by compressing the historical hidden state instead of through the interaction between each element. In contrast, the self-attention of the Transformer is invariant to the permutation of the input elements. Given the three-dimensional nature (e.g., unstructured and disordered) of point clouds, using a single scanning process often struggles to concurrently capture dependency information across various directions, which makes it difficult to construct global modeling for the RNN-like modes (e.g., Mamba). Thus, in this paper, we introduce a simple yet effective Point Cloud State Space Model (denoted as PointMamba [Uncaptioned image]) with global modeling and linear complexity. Specifically, to enable Mamba to capture the point cloud structure causally, we first use a point tokenizer to generate two types of point tokens via a point scanning strategy, employing two space-filling curves to scan key points from different directions. The first type of token has local modeling capabilities through sequential encoding, with the latest token holding global sequence information. Consequently, the second type of token can achieve global modeling by containing global information that comes from the first type. To make the model as simple as possible, PointMamba only employs plain and non-hierarchical Mamba as the backbone to extract features for given serialized point tokens without bells and whistles. In addition, we demonstrate that PointMamba is very flexible in the pre-training paradigm, where we customize an MAE-like pertaining strategy to provide a good prior, which chooses a random serialization strategy from a pre-defined serialization bank to perform mask modeling.
@@
It should be noted that this paper does not claim algorithmic novelty but rather presents a simple and solid Mamba-based baseline with superior performance for point cloud analysis. Despite no elaborate or complex designs in the model, our PointMamba achieves superior performance on various point cloud analysis datasets (Fig. 1(a)). Besides the superior performance, thanks to the linear complexity of Mamba, we show the surprisingly low computational cost [^1] Note that we removed the tokenizer of both Point-MAE and PointMamba , directly fed with a predefined sequence, to better illustrate the structural efficiency., as shown in Fig. 1(b)-(c). These notable results underscore the potential of SSM in 3D vision-related tasks.
@@ 2.1 Point Cloud Transformers
Standard Transformers can be smoothly integrated into autoencoders using an encoder-decoder design he2022masked , which makes this structure ideal for pre-training and leads to significant performance improvements in downstream point cloud analysis tasks yu2022point ; pang2022masked ; chen2023pointgpt . However, the attention mechanism has a time complexity of $O(n^{2}d)$, where $n$ represents the length of the input token sequence and $d$ represents the dimension of the Transformer. This implies that as the input sequence grows, the operational efficiency of the Transformer is significantly constrained.
@@ 2.2 State Space Models
To the best of our knowledge, there are limited works that study SSMs for point cloud analysis. In this work, we delve into the potential of Mamba in point cloud analysis and propose PointMamba [Uncaptioned image], which achieves superior performance and significantly reduces computational costs.
@@ 3 Preliminaries
Space-filling curve. Space-filling curves are paths that traverse every point within a higher-dimensional discrete space while maintaining spatial proximity to a certain degree. Mathematically, they can be defined as a bijective function $\Phi:\mathbb{Z}\to\mathbb{Z}^{3}$ for point clouds. Our PointMamba focuses on the Hilbert space-filling curve and a transposed variant (called Trans-Hilbert), known for their effective locality preservation. This implies that data points close to each other in $\mathbb{Z}$ space stay close even after transformation to $\mathbb{Z}^{c}$. We note that some methods wu2023point ; wang2023octformer utilize space-filling curves to partition the point cloud for capturing spatial contexts, whereas our work mainly focuses on transferring the point clouds to serialization-based sequences and combine with Mamba to implement global modeling. The motivation and objective are different.
@@ 4.2 The serialization-based mask modeling
Figure 4: The details of our proposed serialization-based mask modeling. During the pre-training, we randomly choose one space-filling curve to generate the serialized point tokens for mask modeling.
@@ 5.1 Implementation details
SSM is new to 3D point cloud analysis, with no existing works detailing the specific implementation. To handle different resolutions of the input point cloud, we divide them into different numbers of patches with a linear scaling (e.g., $M=1024$ input points are divided into $n=64$ point patches), with each patch containing $k=32$ points determined by the KNN algorithm. The PointMamba encoder has $N=12$ vanilla Mamba blocks, each Mamba block featuring $C=384$ hidden dimensions. For the pre-training, we utilize ShapeNetCore chang2015shapenet as the dataset, following previous methods yu2022point ; pang2022masked ; chen2023pointgpt . In addition, we utilize 4 $\times$ Mamba blocks as the decoder to reconstruct the masked point clouds. More details can be found in the Appendix C.1.
Table 1: Object classification on ScanObjectNN uy2019revisiting . We evaluate PointMamba on three variants, with PB-T50-RS being the most challenging. Overall accuracy (%) is reported. indicates that using simple rotational augmentation dong2022autoencoders for training.
@@ 5.2 Compared with Transformer-based counterparts
This paper aims to explore the potential of replacing the Transformer with Mamba for point cloud analysis. Thus, in the following experiments, we mainly compare with the state-of-the-art vanilla Transformer-based point cloud analysis methods.
@@
5.3 Analysis and ablation study
The effect of Selective SSM. The key of S6 models or Mamba gu2023mamba is the SSM with the selective mechanism. We prove that as a unidirectional modeling method, SSM can be analogous to masked self-attention, ensuring each position can only attend to previous positions (the detailed proof can be found in the Appendix A.1). Thus, as shown in Tab. 10, we analyze the effect of selective SSM by removing it (i.e., identity setting) or replace with masked self-attention or MLP (an illustration is shown in Fig. 12 from Appendix). Compared with the identity setting, the selective SSM brings notable improvement, indicating the effectiveness of introducing global modeling from SSM. Note that while a very recent method, MambaOut yu2024mambaout , thinks the SSM of Mamba might negatively impact image classification tasks, our findings demonstrate that this is not the case for point cloud analysis tasks. Another interesting thing is that when Selective SSM is replaced with masked self-attention, the performance is even lower than that of the identity setting. We argue the main reason is that masked self-attention is hard to combine with the Gated MLP shazeer2020glu used in default Mamba, leading to optimized difficulty, which might need to be explored in the future.
@@ Appendix B Concurrent Related Works
Some recent works on state space models for point cloud analysis appeared before the submission deadline. This section discusses the differences between these methods and our PointMamba.
@@
Mamba3D han2024mamba3d is another recently proposed method. To obtain better global features, Mamba3D introduces an enhanced Vision Mamba zhu2024vision block, which includes both a token forward SSM and a backward SSM that operates on the feature channel. And it proposes a Local Norm Pooling block to extract local geometric features.
As shown in Tab. 3, our PointMamba surpasses these concurrent methods by offering superior performance with reduced computational overhead. Note that our method is extremely simple and without complex design, which utilizes the vanilla Mamba block and abstains from incorporating modular designs from other baselines, thereby maintaining simplicity and efficiency in our approach. We believe such method can better illustrate the potential of SSM in point cloud analysis tasks.
@@ C.1 Implement details
Downstream tasks Details. Fig 2 shows the pipeline of PointMamba for classification tasks. We report the overall accuracy without voting on the challenging ScanObjectNN uy2019revisiting using 2,048 input points, and on ModelNet40 wu20153d using 1,024 input points. For segmentation on ShapeNetPart yi2016scalable , as shown in Fig. 11, we use random, Hilbert, and Trans-Hilbert serializations, with order indicators applied on Hilbert/Trans-Hilbert serializations. Features from the 3-rd, 7-th, and last layer are pooled as global features after a simple feature fusion. These global features are then concatenated with per-point features and sent to the segmentation head. More detail can be found in Tab. 4.
The different variant of PointMamba. In Sec. 5.3, we analyze the influence of different variant of PointMamba to demonstrate the effectiveness of using select SSM mechanism. We provide an intuitive illustration to better understand these variants, as shown in Fig. 12.
@@ D.1 Mask modeling visualization
As in Sec. 4.2, we customize a simple yet effective serialization-based mask modeling paradigm. By randomly masking about 60% of serialization-based point tokens, an asymmetric vanilla Mamba autoencoder is utilized to extract the point feature, with a simple prediction head for reconstruction. In Fig 13, we present qualitative results of mask modeling on ShapeNet validation set. Despite a 60% masking ratio, our PointMamba effectively reconstructs the masked patches, providing a strong self-supervised knowledge for downstream tasks.

### v5

@@ Abstract
Transformers have become one of the foundational architectures in point cloud analysis tasks due to their excellent global modeling ability. However, the attention mechanism has quadratic complexity, making the design of a linear complexity method with global modeling appealing. In this paper, we propose PointMamba [Uncaptioned image], transferring the success of Mamba, a recent representative state space model (SSM), from NLP to point cloud analysis tasks. Unlike traditional Transformers, PointMamba employs a linear complexity algorithm, presenting global modeling capacity while significantly reducing computational costs. Specifically, our method leverages space-filling curves for effective point tokenization and adopts an extremely simple, non-hierarchical Mamba encoder as the backbone. Comprehensive evaluations demonstrate that PointMamba achieves superior performance across multiple datasets while significantly reducing GPU memory usage and FLOPs. This work underscores the potential of SSMs in 3D vision-related tasks and presents a simple yet effective Mamba-based baseline for future research. The code will be made available at https://github.com/LMD0311/PointMamba.
@@ 1 Introduction
Point cloud analysis is one of the fundamental tasks in computer vision and has a wide range of real-world applications shi2020points ; wu2023point ; chen2023voxelnext , including robotics, autonomous driving, and augmented reality. It is a challenging task due to the intrinsic irregularity and sparsity of point clouds. To address the issues, there has been rapid progress in deep learning-based methods pang2022masked ; liu2022masked ; qi2017pointnet++ ; qian2022pointnext , consistently pushing the performance to the new record.
@@
Inspired by this, this paper aims to unlock the potential of SSM in point cloud analysis tasks, discussing whether it can be a viable alternative to Transformers in this domain. Through a series of pilot experiments, we find that directly using the pioneering SSM, Mamba gu2023mamba , can not achieve ideal performance. We argue that the main inherent limitation comes from the unidirectional modeling employed by the default Mamba, as the context is obtained by compressing the historical hidden state instead of through the interaction between each element. In contrast, the self-attention of the Transformer is invariant to the permutation of the input elements. Given the three-dimensional nature (e.g., unstructured and disordered) of point clouds, using a single scanning process often struggles to concurrently capture dependency information across various directions, which makes it difficult to construct global modeling for the RNN-like modes (e.g., Mamba).
Therefore, we introduce a simple yet effective Point Cloud State Space Model (denoted as PointMamba [Uncaptioned image]) with global modeling and linear complexity. Specifically, to enable Mamba to capture the point cloud structure causally, we first use a point tokenizer to generate two types of point tokens via a point scanning strategy, employing two space-filling curves to scan key points from different directions. As a result, the unstructured 3D point clouds can be transformed into a regular sequence. The first type of token has local modeling capabilities through sequential encoding, with the latest token holding global sequence information. Consequently, the second type of token can achieve global modeling by containing global information that comes from the first type. Besides, we propose an extremely simple order indicator to maintain the distinct spatial characteristics of different scanning when training, preserving the integrity of the spatial information.
To make the model as simple as possible, PointMamba only employs plain and non-hierarchical Mamba as the backbone to extract features for given serialized point tokens without bells and whistles. We demonstrate that PointMamba is very flexible in the pre-training paradigm, where we customize an MAE-like pertaining strategy to provide a good prior, which chooses a random serialization strategy from a pre-defined serialization bank to perform mask modeling. It facilitates the model to exact the general local relationships from different scanning perspectives, better matching the requirement of indirection modeling of mamba.
@@
Despite no elaborate or complex designs in the model, our PointMamba achieves superior performance on various point cloud analysis datasets (Fig. 1(a)). Besides the superior performance, thanks to the linear complexity of Mamba, we show the surprisingly low computational cost [^1] Note that we removed the tokenizer of both Point-MAE and PointMamba , directly fed with a predefined sequence, to better illustrate the structural efficiency., as shown in Fig. 1(b)-(c). These notable results underscore the potential of SSM in 3D vision-related tasks.
@@ 2.1 Point Cloud Transformers
Standard Transformers can be smoothly integrated into autoencoders using an encoder-decoder design he2022masked , which makes this structure ideal for pre-training and leads to significant performance improvements in downstream point cloud analysis tasks yu2022point ; pang2022masked ; chen2023pointgpt ; zha2024lcm ; zha2023instance . However, the attention mechanism has a time complexity of $O(n^{2}d)$, where $n$ represents the length of the input token sequence and $d$ represents the dimension of the Transformer. This implies that as the input sequence grows, the operational efficiency of the Transformer is significantly constrained.
@@ 2.2 State Space Models
To the best of our knowledge, there are limited works that study SSMs for point cloud analysis. In this work, we delve into the potential of Mamba in point cloud analysis and propose PointMamba, which achieves superior performance and significantly reduces computational costs.
@@ 3 Preliminaries
Space-filling curve. Space-filling curves are paths that traverse every point within a higher-dimensional discrete space while maintaining spatial proximity to a certain degree. Mathematically, they can be defined as a bijective function $\Phi:\mathbb{Z}\to\mathbb{Z}^{3}$ for point clouds. Our PointMamba focuses on the Hilbert space-filling curve hilbert1935stetige and its transposed variant (called Trans-Hilbert), both of which are recognized for effectively preserving locality, ensuring that data points close in $\mathbb{Z}$ space remain close after transformation to $\mathbb{Z}^{3}$. We note that some methods wu2023point ; wang2023octformer utilize space-filling curves to partition the point cloud for capturing spatial contexts, whereas our work mainly focuses on transferring the point clouds to serialization-based sequences and combine with Mamba to implement global modeling. The motivation and objective are different.
@@ 4.2 The serialization-based mask modeling
Figure 4: The details of our proposed serialization-based mask modeling. During the pre-training, we randomly choose one space-filling curve to generate the serialized point tokens for mask modeling, and different serialized point tokens have different order indicators.
@@ 5.1 Implementation details
SSM is new to 3D point cloud analysis, with no existing works detailing the specific implementation. To handle different resolutions of the input point cloud, we divide them into different numbers of patches with a linear scaling (e.g., $M=1024$ input points are divided into $n=64$ point patches), with each patch containing $k=32$ points determined by the KNN algorithm. The PointMamba encoder has $N=12$ vanilla Mamba blocks, each Mamba block featuring $C=384$ hidden dimensions. For the pre-training, we utilize ShapeNetCore chang2015shapenet as the dataset, following previous methods yu2022point ; pang2022masked ; chen2023pointgpt . In addition, we utilize 4 $\times$ Mamba blocks as the decoder to reconstruct the masked point clouds.
Table 1: Object classification on the ScanObjectNN dataset uy2019revisiting . We evaluate PointMamba on three variants, with PB-T50-RS being the most challenging. Overall accuracy (%) is reported. Param. denotes the number of tunable parameters during training. indicates that using simple rotational augmentation dong2022autoencoders for training.
@@
Point-MAE+IDPT zha2023instance ICCV 23 Transformer 1.7 7.2 91.22 90.02 84.94
Point-MAE+DAPT zhou2024dynamic CVPR 24 Transformer 1.1 5.0 90.88 90.19 85.08
@@ 5.2 Compared with Transformer-based counterparts
This paper aims to unlock the potential of Mamba in point cloud tasks, discussing whether it can be a viable alternative to Transformers. Thus, in the following experiments, we mainly compare with the state-of-the-art vanilla Transformer-based point cloud analysis methods.
@@
5.3 Analysis and ablation study
The effect of Selective SSM. The key of S6 models or Mamba gu2023mamba is the SSM with the selective mechanism. We prove that as a unidirectional modeling method, SSM can be analogous to masked self-attention, ensuring each position can only attend to previous positions (the detailed proof can be found in the Appendix A.1). Thus, as shown in Tab. 11, we analyze the effect of selective SSM by removing it (i.e., identity setting) or replace with masked self-attention or MLP (an illustration is shown in Fig. 7). Compared with the identity setting, the selective SSM brings notable improvement, indicating the effectiveness of introducing global modeling from SSM. Note that while a very recent method, MambaOut yu2024mambaout , thinks the SSM of Mamba might negatively impact image classification tasks, our findings demonstrate that this is not the case for point cloud analysis tasks. Another interesting thing is that when Selective SSM is replaced with masked self-attention, the performance is even lower than that of the identity setting. We argue the main reason is that masked self-attention is hard to combine with the Gated MLP shazeer2020glu used in default Mamba, leading to optimized difficulty, which might need to be explored in the future.
@@ Appendix B Concurrent Related Works
Some works on state space models for point cloud analysis appeared recently. This section discusses the differences between these methods and our PointMamba.
@@
Mamba3D han2024mamba3d is another recently proposed method. To obtain better global features, Mamba3D introduces an enhanced Vision Mamba zhu2024vision block, which includes both a token forward SSM and a backward SSM that operates on the feature channel. It proposes a Local Norm Pooling block to extract local geometric features.
PointTramba wang2024pointramba introduces a hybrid approach that integrates Transformers and Mamba. It segments point clouds into groups and utilizes Transformers to capture intra-group dependencies, while Mamba models inter-group relationships using a bi-directional, importance-aware ordering strategy.
As shown in Tab. 3, our PointMamba surpasses these concurrent methods by offering superior performance with reduced computational overhead. Note that our method is extremely simple and without complex design, which utilizes the vanilla Mamba block and abstains from incorporating modular designs from other baselines, thereby maintaining simplicity and efficiency in our approach. We believe such a method can better illustrate the potential of SSM in point cloud analysis tasks.
@@ C.1 Implement details
Downstream tasks Details. Fig. 2 shows the pipeline of PointMamba for classification tasks. We report the overall accuracy without voting on the challenging ScanObjectNN uy2019revisiting using 2,048 input points, and on ModelNet40 wu20153d using 1,024 input points. For segmentation on ShapeNetPart yi2016scalable , as shown in Fig. 12, we use random, Hilbert, and Trans-Hilbert serializations, with order indicators applied on Hilbert/Trans-Hilbert serializations. Features from the 3-rd, 7-th, and last layer are pooled as global features after a simple feature fusion. These global features are then concatenated with per-point features and sent to the segmentation head. More detail can be found in Tab. 4.
C.2 Additional ablation study
Figure 13: The effect of masking strategy. The pre-training loss (× 1000) along with fine-tuning accuracy (%) are reported.
Masking ratio Loss OBJ-BG OBJ-ONLY
0.4 2.01 92.60 90.70
0.6 1.97 94.32 92.60
0.8 2.33 93.46 90.17
0.9 2.00 92.43 91.05
Figure 14: The effect of classification token. Fine-tuning accuracy (%) are reported.
Methods OBJ-BG OBJ-ONLY
Before the sequence 93.63 91.05
After the sequence 94.32 90.19
Middle the sequence 93.98 90.71
Max Pool 93.39 91.36
Average Pool 94.32 92.60
In this section, we do additional ablation studies on several hyper-parameters.
Masking strategy for pre-training. By employing a serialization-based mask modeling paradigm, our PointMamba achieves superior performance. To find a proper masking strategy for our method, we compare two types of masking with varying ratios. The block masking yu2022point masks geometrically proximate point cloud patches, leading to a more challenging reconstruction target. In Tab. 14, we experimentally find that masking 60% of point patches by randomly choosing can achieve good performance.
Usage of classification token. Previous works dosovitskiy2020image ; yu2022point ; pang2022masked often use a classification token $[\mathtt{CLS}]$ as a global token for classification. As in Tab. 14, we find that without $[\mathtt{CLS}]$ and utilizing only the average pooling of the final block’s output yields the best results for PointMamba.
@@ D.1 Mask modeling visualization
As in Sec. 4.2, we customize a simple yet effective serialization-based mask modeling paradigm. By randomly masking about 60% of serialization-based point tokens, an asymmetric vanilla Mamba autoencoder is utilized to extract the point feature, with a simple prediction head for reconstruction. In Fig. 15, we present qualitative results of mask modeling on ShapeNet validation set. Despite a 60% masking ratio, our PointMamba effectively reconstructs the masked patches, providing a strong self-supervised knowledge for downstream tasks.